x-y+9z=-27
2x-4y-z=-1
3x+6y-3z=27
I have tried and tried on this problem for an hour and have gotten different answers each time. All I need to know is what are the values of x, y, and z?
Thank you.

Answer by scott8148(6628)

multiplying 1st eqn by -2 and adding to 2nd eqn ___ -2y - 19z = 53 (A)

multiplying 1st eqn by -3 and adding to 3rd eqn ___ 9y - 30z = 108 (B)

multiplying (A) by 9, and (B) by 2 and adding ___ -231z = 693

solve for z and then substitute back to find y and x